OpenStudy (anonymous):

Flying against the wind,an airplane travels 2,880 miles in 4 1/2 hours. flying with the wind,the airplane can travel the same distance 4hours .find the speed of the plane in calm air A)680 B)40mph c)50mph D)610mph

OpenStudy (anonymous):

Let s and w be the speed of the plane and the wind respectively. Solve the following for s and w:\[\left\{\frac{9}{2}(s-w)\text{ = }2880,4(s+w)\text{ = }2880\right\}\]\[{s = 680 mph, w = 40 mph} \]

OpenStudy (anonymous):

Thank you